OpenStudy (anonymous):

Given the function g(x) = -4x + 5, find g(-4).

OpenStudy (bahrom7893):

just plug in -4 instead of x: -4*(-4) + 5 = 16+5 = 21

OpenStudy (rulnick):

bahrom is right

OpenStudy (bahrom7893):

yea i tend to be right most of the time lol :)

OpenStudy (bahrom7893):

thanks rulnick :)

OpenStudy (anonymous):

Ohk thanks you guys I really appreciate it :)

OpenStudy (rulnick):

welcomes